(a)Consider the binomial distribution with parameters.n, \piDraw 4 figures that will show qualitatively how m_{x}, \sigma ^2_{x} vary with n, \pi

suppose we toss n = 100 unfair coins, with an unknown \pi

(b) what is the expected number of heads out of n? (the answer depends on \pi )

(c) what is the typical deviation of the number of heads out of n? (the answer depends on \pi )

(d) what is the largest typical deviation of the number of heads out of n? (the answer is a number) Hint: consult your graph of \sigma^2 vs \pi in part a)
Answer #1

a)

\mu_x = n \pi

first and second graph will increase linearly keeping the other parameter constant

\sigma ^2_x = n \pi(1-\pi )

third graph also linear with respect to n

fourth graph will be inverted parabola , maximum at pi = 1/2 and 0 when pi = 0 and 1

b)

\mu_x = n \pi

c)

\sigma ^2_x = n \pi(1-\pi )

\sigma_x =sqrt(n \pi (1-\pi))

d)

\sigma ^2_x = n \pi(1-\pi )

largest value when \pi = 1/2

\sigma_x =sqrt(n/4)

Please rate